### 12 Dec 2023 v2

Abstract
Multilevel regression and poststratification (MRP) has been used extensively to adjust convenience and low-response surveys to make population and subpopulation estimates. For this method, model validation is particularly important, but recent work has suggested that simple aggregation of individual prediction errors does not give a good measure of the error of the population estimate. In this manuscript we provide a clear explanation for why this occurs, propose two scoring metrics designed specifically for this problem, and demonstrate their use in three different ways. We demonstrate that these scoring metrics correctly order models when compared to true goodness of estimate, although they do underestimate the magnitude of the score.
Keywords: sample surveys, model validation, Bayesian statistics
1 Introduction
Surveys are increasingly reliant on model-based estimates to overcome low response rates and non-probability samples. In this manuscript we focus on a model-based technique, multilevel regression and poststratification (MRP; Gelman & Little, 1997, Park et al., 2006 and Lax & Phillips, 2009b). This method has been widely used in political science (e.g., Lax & Phillips, 2009a, Ghitza & Gelman, 2013, and Wang et al., 2015) and increasingly deployed in social (e.g., Alsalti et al., 2023, and Kolczynska et al., 2020) and health sciences (e.g., Machalek et al., 2022, and Downes et al., 2018). As with all methods of survey adjustment, the method relies on assumptions of model suitability.
MRP has three stages (see, e.g., Lopez-Martin et al., 2022). First, a model (originally a Bayesian multilevel model, but increasingly other regularising models, e.g., Ornstein, 2020, Bisbee, 2019, and Gao et al., 2021) is fit to the observed sample data conditional on discrete predictors (typically geographic and demographic) whose joint distribution in the population is known or can be estimated. Second, a poststratification table is constructed to represent this population joint distribution, ideally using census data or high quality official statistics. The model from the sample is then used to predict the expected value for each cell defined by $X$ in the population. Lastly, the population totals in each cell are used to weight the cell estimates to obtain an estimate for the population (or subpopulation). The three steps can be remembered as model-predict-aggregate.
The goodness of the population estimate will be heavily reliant on the goodness of the model predictions. This means that it is essential to consider how to fairly validate the model. Given the emphasis on prediction in this method, previous work investigated the efficacy of the sum of leave-one-out(LOO) expected log predictive density (ELPD) in estimating true error (Kuh et al., 2022). This is a commonly used metric for model goodness in the Bayesian literature, and is easily implemented through the LOO package (Vehtari et al., 2021). Kuh et al. (2022) also considered a weighted alternative to this metric (initially proposed by Lumley & Scott, 2015) to account for the non-random sample.
Little & Vartivarian (2005) demonstrate that the required adjustment variables to produce an unbiased estimate are those that are related to both the probability of inclusion and the probability of the outcome (which Kuh et al., 2022 refer to as bias variables). However, neither standard sum of LOO ELPD nor the weighted alternative correctly distinguish between these bias variables and precision variables (variables with are strongly related to the model outcome but not sample inclusion). This is a particular challenge for MRP, as each variable in the model needs to be found or estimated in the population in order to create a poststratification matrix. Kuh et al. (2022) also note additional challenges in this space, such as the difference in preferred model for different subpopulations.
1.1 Individual scores and aggregate scores
We argue that the underlying difficulty described in Kuh et al. (2022) is that the best model for an average over individuals may not correspond to the best model for an aggregate estimate. To understand why, consider a simple example of a sample of size 2, where these individuals have a true outcome value 0. We will say $y_{\mathrm{true}}=\{0,0\}$. The mean of $y_{true}$ is $E(y_{\mathrm{true}})=\frac{0+0}{2}=0$. Now we have two potential models to estimate $y_{\mathrm{true}}$. Each produces an estimate (column 2, Table 1), from which we can calculate the squared error and the mean squared error over both individuals. We can also use these estimates to create an estimate of the mean, and from this calculate the squared error of the mean. As we can see, while model 1 is preferred by the mean squared individual error, model 2 is far preferred at estimating the mean.
Individual Mean
$\hat{y_{i}}$ $(\hat{y_{i}}-y_{i})^{2}$ $\frac{1}{2}\sum(\hat{y_{i}}-y_{i})^{2}$ $\frac{1}{2}\sum\hat{y_{i}}$ $(\frac{1}{2}\sum\hat{y_{i}}-\frac{1}{2}\sum y_{i})^{2}$
Model 1 {$0,1$} {$0,1$} 0.5 0.5 0.25
Model 2 {$-2,2$} {$-2,2$} 4 0 0
Table 1: Demonstration of difference between mean of individual squared error and squared error of mean in a simple example of $n=2$. True outcome $y_{i}=0$.
This simple example demonstrates what we believe to be one of the underlying challenges of validating multilevel regression and poststratification (MRP) models. An MRP based estimator is by definition an estimator that uses cellwise estimates to compose a population or subpopulation estimates. Previously we had investigated using sum of individual goodness to estimate the goodness of the MRP estimate, which is not effective (Kuh et al., 2022).
1.2 Notation and key terms
Consider a finite population $P$ with $N$ individuals, where interest lies in a binary variable $Y$. Whilst some authors have used a continuous outcome variable (Liu et al., 2023), the prevalent applications currently focus on estimating probabilities. Assume also a set of $k$ predictors whose values are known in the population and are represented by a $N\times k$ matrix, $\mathbf{X}$. Assume the predictors are discrete, with the $k^{th}$ variable, $X^{(k)}$, having $l^{(k)}$ levels. This means the distribution of $X$ in the population can be represented by the number of items within each of the $J=\prod_{k}l^{(k)}$ possible configurations of predictors.
The population mean of variable of interest can be written as,
$\operatorname{E}_{\mathrm{P}}(Y)=\frac{\sum_{i=1}^{N}Y_{i}}{N}.=\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}=1)}{N},$ (1)
where $\mbox{Pr}(Y_{i\in j}=1)=\sum_{i\in j}Y_{i}/N_{j}$.
Next take a sample of size $n$. The observations in this sample are assumed to be random conditional on the predictors, $X$. The probability of inclusion in the sample for every $j^{th}$ cell is denoted $\pi_{j}$.
Following traditional sampling notation, we use lower case for the sample and upper case for the population, and where a matrix is needed, it will be presented in boldface. Our goal is to estimate the population mean by first estimating the probability of the outcome in a each cell in the population using the sample. We denote this quantity for the $j^{t}h$ cell as $\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\theta)$. We then use this to estimate the population mean, $E_{P}(Y|y,\textbf{x},\theta)$. We do this by modifying (1) into
$\operatorname{E}_{\mathrm{P}}(Y|y,\textbf{x},\theta)=\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\theta)}{\sum_{j=1}^{J}N_{j}}.$ (2)
Our goal is to identify the best model $M$ and estimate the parameters $\theta$, denoted $\hat{\theta}$, using the sample. We can then use this to estimate $\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})$, which can then be used in (2).
$\operatorname{E}_{\mathrm{P}}(Y|y,\textbf{x},\hat{\theta})=\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})}{\sum_{j=1}^{J}N_{j}}.$ (3)
In this paper we discuss aggregation at a range of different levels. For clarity we describe all terms with a simple example. Consider a survey of voting intention. The smallest unit is the individual, the person who is being surveyed. These individuals can be aggregated in different ways. For example, we could aggregate all individuals with a specific set of demographic qualities (e.g., aged 18–25, female, college educated). We call this a cell if these three demographics are the full set of $X$ variables. We could also aggregate at a particular level (e.g., 18-25 year) of a variable (age). This is a special case of a subpopulation (also known as a small area), which is where we estimate part of the population. In the limit the subpopulation is the full population (e.g., all US voters).
1.3 Contributions
In this manuscript we
  • Propose an alternative scoring metric when estimating the probability of a binary outcome in the population (Section 2)
  • Demonstrate the use of this method in a leave-one-cell-out cross validation scheme, and demonstrate that unlike previous work, this method retains the correct model ordering (Section 3)
  • Apply this method to subpopulation estimation (Section 4)
  • Extend the leave-one-cell-out cross validation to a leave-one-cell-out reference validation score where instead of using the sample as a proxy for the cellwise population truth we use a reference model estimate (Section 5)
  • The leave-one-cell-out approach requires all cells to be observed in the sample. We also propose an approximate leave-one-cell-out reference approach for use when not all cells are observed in the sample (Section 5), and a combined validation approach that incorporates leave-one-cell-out cross validation and leave-one-cell-out reference validation (Section 5).
Together we intend for these contributions to provide a convincing argument that the proposed scoring metric is more appropriate for scoring models for MRP. The structure of this paper is as follows. In Section 2 we describe our scoring rule adjustment. In Section 3 we apply our scoring rules in a leave-one-cell-out cross validation scheme. In Section 4 we apply our scoring rules to score subpopulation estimates using approximate leave-one-cell-out cross validation. In Section 5 we discuss the use of reference validation scores.
2 Aggregate scoring from cellwise error
In this section we demonstrate that by decomposing a score of the MRP population mean into its cellwise components, we can produce an estimate of the population mean score using cellwise goodness. We complete this for a squared error score ($\operatorname{Error}^{2}$) for the population mean and the continuous ranked probability score (CRPS). This decomposition may be possible for other scoring rules.
2.1 Squared Error
First consider the squared error for the population mean,
$\operatorname{Error}_{\mathrm{MRP}}^{2}=(E_{P}(Y|\theta,y,\textbf{x})-E_{P}(Y))^{2}.$ (4)
Using (1) and (3), we can instead write:
$\operatorname{Error}_{\mathrm{MRP}}^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})}{\sum_{j=1}^{J}N_{j}}-\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}=1)}{\sum_{j=1}^{J}N_{j}}\Bigg{)}^{2},$ (5)
which can be simplified to
$\operatorname{Error}_{\mathrm{MRP}}^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})\,-\,\mbox{Pr}(Y_{i\in j}=1))}{\sum_{j=1}^{J}N_{j}}\Bigg{)}^{2}.$ (6)
The sum is on the inside of the squared term, which demonstrates why simply summing the squared error of the individuals or cells (as is often done when scoring models; see Gelman et al., 2014) would not provide an acceptable score for the population mean. For clarity, the mean of the squared error of the cells, adjusted for sample representation would be
$\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})-\mbox{Pr}(Y_{i\in j}=1))^{2}}{\sum_{j=1}^{J}N_{j}}.$ (7)
2.2 Continuous ranked probability scores
The continuous ranked probability score (CRPS) is an extension of squared error that incorporates the full probability distribution. This relationship suggests that the CRPS for the population level MRP estimate be similarly decomposed. CRPS is described as
[…]
where $F$ is the CDF of the predictive distribution and $x$ is the true value (Gneiting & Raftery, 2007). One benefit of this score is that it can be implemented through approximate draws of this distribution (Equation (21) of Gneiting & Raftery, 2007). To harmonise notation, we denote $\phi$ as the posterior estimate for the population expectation.
$\operatorname{CRPS}(\phi,E_{P}(Y))=\frac{1}{2}E(\lvert\phi-\phi^{\prime}\rvert)-E(\lvert\phi-E_{P}(Y)\rvert).$ (9)
In our setting we focus on cases where we use posterior draws to estimate expectation. Denoting the $b^{th}$ posterior draw as $\mathbf{\phi}^{b}$ and $\mathbf{\phi}^{\prime}$ as a permutation of the posterior draws of $\mathbf{\phi}$, we can write this as
$\operatorname{CRPS}(\phi,E_{P}(Y))=\frac{1}{2}\frac{1}{B}\sum_{b}^{B}\lvert\phi^{b}-\phi^{\prime b}\rvert-\frac{1}{B}\sum_{b}^{B}\lvert\phi^{b}-E_{P}(Y)\rvert.$
The $\phi$ represents our MRP estimate, and $E_{P}(Y)$ the population mean truth, so we can rewrite this as
$\!\!\!\!\!\!\!\!\operatorname{CRPS}_{\mathrm{MRP}}(Y,y,\hat{\theta},\mathbf{x})=$
$\frac{1}{B}\sum_{b}^{B}\Bigg{(}\frac{1}{2}\ \Bigg{\lvert}\frac{\sum_{j=1}^{J}(\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{b},\textbf{x})*N_{j})}{\sum_{j=1}^{J}N_{j}}\,-\,\frac{\sum_{j=1}^{J}(\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{\prime b},\textbf{x})*N_{j})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\,-$
$\ \ \ \ \ \ \ \ \Bigg{\lvert}\frac{\sum_{j=1}^{J}(\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{b},\textbf{x})*N_{j})}{\sum_{j=1}^{J}N_{j}}\,-\,\frac{\sum_{j=1}^{J}(\mbox{Pr}(Y_{i\in j}=1)*N_{j})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$
Within each summation, this can be reordered to produce a weighted sum of cellwise components:
$\!\!\!\!\operatorname{CRPS}_{\mathrm{MRP}}(Y,y,\hat{\theta},\mathbf{x})=\frac{1}{B}\sum_{b}^{B}\Bigg{(}\frac{1}{2}\ \Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{\prime b},\textbf{x})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\,-\\ \Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}=1)\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$ (10)
As with squared error, the absolute value is outside of the summation, rather than inside as we would expect for a usual sum of the cellwise scores:
$\begin{split}\operatorname{CRPS}(Y,y,\hat{\theta},\mathbf{x})=\frac{1}{N}\sum_{j=1}^{J}N_{j}\Bigg{(}\frac{1}{2}\frac{1}{B}\sum_{b}^{B}\Big{\lvert}\mbox{Pr}(Y_{i\in j}=1|\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j}=1|\hat{\theta}^{\prime b})\Big{\rvert}\,-\,\\ \frac{1}{B}\sum_{b}^{B}\Big{\lvert}\mbox{Pr}(Y_{i\in j}=1|\hat{\theta}^{b})-\mbox{Pr}(Y_{j})\Big{\rvert}\Bigg{)}.\end{split}$ (11)
2.3 Demonstration of equivalence
Here we demonstrate the equivalency of the cellwise decomposition to the original population score. To achieve this, we directly calculate the squared error and CRPS for the MRP estimate at the population level and then also calculate this using the cellwise estimates.
For simplicity, we borrow the simulation scenario in Example 1 of Kuh et al. (2022), but add an additional constraint that all cells in the population must be observed at least once in the sample (Kuh et al., 2022 had a less restrictive constraint that at least one observation every level needed to be observed). In this example, four independent, normally distributed ($\mbox{normal}(\mu=0,\sigma=2)$) variables are created. The probability of the outcome ($\mbox{Pr}(Y_{i}=1)$) and the probability of being included in the sample ($\pi_{i}$) were created using the following relationships:
[…]
We note from this that $X_{2}$ is a precision variable (related strongly to the outcome but weakly to the probability of inclusion in the sample), and $X_{4}$ is a bias variable (related strongly to both). According to Little & Vartivarian (2005), models with $X_{4}$ and not $X_{2}$ should be strongly preferred over models with $X_{2}$ and not $X_{4}$.
As we are focusing on a binary outcome, we create observed binary outcomes $y_{i}$ for every individual in the population using $\mbox{Pr}(y_{i}=1)$. We also discretize the $X^{(k}$ variables into $5$ groups of equal range, which we use to model.
We simulate with a population size of $20,000$, twice the size of Kuh et al. (2022). We sample from the population firstly so that every cell has one observation and the secondly with probability $\pi_{i}$. The sample size is held constant at $1,000$. The population creation and sampling is repeated $100$ times. As the outcome is binary, we choose to fit a binomial likelihood of counts per cell $j$. We also only fit a subset of the models used by Kuh et al. (2022), focusing on the following models:
  • Full model: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{2})}_{j}+\alpha^{(X_{3})}_{j}+\alpha^{(X_{4})}_{j}),$
  • Precision variable model: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{2})}_{j}+\alpha^{(X_{3})}_{j}),$
  • Bias variable model: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{3})}_{j}+\alpha^{(X_{4})}_{j}),$
  • Nuisance variable model: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}+\alpha^{(X_{3})}_{j}),$
where $\alpha^{k}_{j}\sim\operatorname{normal}(0,\sigma^{(k)}$), $\sigma^{(k)}\sim t(3,0,2.5)$ and $\beta_{0}\sim t(3,0,2.5)$, which are the standard priors in the brms R package (Bürkner, 2017, 2018).
Figure 1: Comparison of a direct score of the MRP estimate (using known population truth, $y$-axis) against an indirect score of the MRP estimate (using the known population truth for each cell, $x$-axis). Colour of point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents an identity line. As the points all fall either directly or with a small amount of variation, we find evidence that this decomposition of the scoring rule is valid.
For each sample iteration and model we perform the following steps:
  1. 1.
    Create a population.
  2. 2.
    Create a sample from this population so that each cell has at least one individual observed, and then according to $\pi_{i}$.
  3. 3.
    Fit all four models on this sample using brms with default settings.
  4. 4.
    For each model we calculate
    1. (a)
      The population mean estimate using MRP and score it compared to the truth using CRPS and squared error,
    2. (b)
      An estimate for each cell and compare to the true $\mbox{Pr}(Y_{j}=1)$ to calculate the cellwise CRPS and squared error as per (10) and (6).
Figure 1 plots the score for our population mean estimated with MRP against the score calculated through the cellwise error. Up to a small amount of noise, these two scores are equivalent. This provides confirmation for our proposal that the score for a population mean estimate can be decomposed into cell estimates.
3 Estimating MRP goodness using sample goodness
3.1 Sample as proxy
Continuing with the previous example, we aim to score our model without population truth using the observed values in each cell. We first consider a scenario where at least one individual in every cell is observed. We continue this assumption in Section 4 where we consider subpopulation estimates. In Section 4 we relax this assumption to consider techniques when not all cells are observed in the sample.
The results from Section 2 are a useful proof of the equivalence of these two formulations, but it is not directly clear how they might be used in a real world setting where the truth is not known. One example is to implement an approximation of (6) and (10), so that instead of using the population truth we instead approximate the population truth using a cellwise sample estimate which we denote $\mbox{Pr}(y_{i\in j}=1|n_{j})$. In these adaptions, we replace the population truth with this value:
$\widehat{\operatorname{Error}}_{\mathrm{MRP}}^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})-\mbox{Pr}(y_{i\in j}=1|n_{j},y_{j}))}{\sum_{j=1}^{J}N_{j}}\Bigg{)}^{2},$ (14)
$\widehat{\operatorname{CRPS}}_{\mathrm{MRP}}(y,n,\hat{\theta},\mathbf{x})=\frac{1}{2}\frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{\prime b},\textbf{x})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}-\\ \frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|n_{j},y_{j})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$ (15)
When there is only one observation per cell, this probability estimate is either 00 or $1$ accordingly. Figure 2 shows this new estimate on the $x$-axis and the true score on the $y$-axis. In both plots, we see two things relative to Figure 1. Our estimates of the model score are noisier but retain the basic ordering of models previously observed—namely, the separation between models that contain the bias variable as a predictor and models that contain only the precision variable as a predictor. This is important as the standard sum of individual error does not maintain that separation (e.g., see Kuh et al., 2022 for sum of ELPD, and supplementary materials for mean of CRPS and squared error scores). However, the error as estimated in the sample is a considerable underestimate relative to the truth. This could be explained by using the sample without correcting for overfitting.
Figure 2: Comparison of a direct score of the MRP estimate (using known population truth, $y$-axis) against an indirect score (using the known sample truth for each cell, $x$-axis). Colour of point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents the identity. The CRPS scores estimated with the sample are overall higher than the true score (where lower indicates a worse model) and the squared error scores estimated with the sample are overall lower than the true score (where higher indicates a worse model). This indicates that using the sample as a proxy for population truth underestimates the model error.
3.2 Brute-force cross validation
To resolve the underestimation of the magnitude of scores seen in the previous section, we consider a cross validation approach. There is a wide and established literature on cross validation for estimating prediction error from a sample. There are three potential schemes that could be used for this problem:
  1. 1.
    $K$-fold cross validation: In this approach we take split the sample into folds. Then we leave the $k^{th}$ fold out to fit the model and estimate error using that fold. In the context of MRP, we would need to adjust the error estimated in this fold to population. This would require that at least $k$ observations present in every poststratification matrix cell (one for each fold). While this might seem plausible, the size and complexity of survey data means that this is unlikely, even if MRP itself doesn’t require that every cell in the poststratification matrix is observed.
  2. 2.
    Leave-one-cell-out cross validation: In this implementation we leave out one cell of the poststratification matrix, predict the probability using a model fit with the remaining data, and then estimate model error using an estimate of $\mbox{Pr}(y_{j}=1|n_{j})$. In this scenario, only one observation is required per cell because it can be composed using our proposed formulae.
  3. 3.
    Leave-one-level-out (LOLO) cross validation: In this implementation we leave out one entire level of a adjustment variable ($x^{(k)}$), fit the model without it and predict for all observations in that level (this is also referred to as leave-one-group-out in the cross validation literature). However, it is not clear which variable to focus on when multiple variables are in the model. This approach could be useful approach for subpopulation estimation if the subpopulation represents different levels of a variable; see Section 4.
From this set of potential approaches, we feel that option 2 best aligns with the decomposition of total MRP score into cellwise components. This is not simply a special case of a $K$-fold because the individuals scores in each cell are not summed, but instead applied in adaptions to our previous formulae. Specifically we propose replacing our estimate $\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})$ with $\mbox{Pr}(Y_{i\in j}=1|y_{i\notin j},\textbf{x},\hat{\theta})$, the probability of $y=1$ in that cell given that it was not used when the model was fit:
$\widehat{\operatorname{Error}}_{\mathrm{MRP,LOCO}}^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}=1|y_{i\notin j},\textbf{x},\hat{\theta})-\mbox{Pr}(y_{i\in j}=1|n_{j},y_{j}))}{\sum_{j=1}^{J}N_{j}}\Bigg{)}^{2},$ (16)
$\widehat{\operatorname{CRPS}}_{\mathrm{MRP,LOCO}}(y,n,\hat{\theta},\mathbf{x})=\frac{1}{2}\frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{\prime b},\textbf{x})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}-\\ \frac{1}{B}\sum_{b}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|n_{j},y_{j})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$ (17)
This brute-force implementation of leave-one-cell-out (LOCO) is not particularly efficient in terms of time/computational resources, but it does allow us to directly investigate whether the underestimation of error present found previously can be eliminated with this approach. Figure 3 shows that the error estimated with a brute-force LOCO approach is indeed slightly larger than that estimated using the full sample to model and estimate error. However, this is still a considerable underestimate for the true error, as shown by Figure 4.
Figure 3: Comparison of the brute-force leave-one-cell-out approximation (LOCO, $y$-axis) against the the indirect score of the MRP estimate (using the known sample truth for each cell, $x$-axis). Colour of point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents an identity line. The CRPS scores estimated brute force LOCO are overall lower than the true score (where lower indicates a worse model) and the squared error scores estimated with brute force LOCO are overall higher than the true score (where higher indicates a worse model). This indicates that using brute force LOCO at least partially corrects the score underestimate when using the sample sample as a proxy for population truth.
Figure 4: Comparison of a direct score of the MRP estimate (using known population truth, $y$-axis) against an indirect score of the MRP estimate (using a brute-force leave-one-cell-out approach, $x$-axis). Colour of point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents an identity line. The CRPS scores estimated with LOCO are overall higher than the true score (where lower indicates a worse model) and the squared error scores estimated with LOCO are overall lower than the true score (where higher indicates a worse model). This indicates that using LOCO as a proxy for population truth still underestimates the model error.
Despite the underestimate of the magnitude of the error, we believe that this approach is still useful to practitioners who wish to evaluate the relative efficacy of different models when estimating the population mean. Unlike previous attempts, ordering is maintained. We advocate that the magnitude of the error be interpreted with caution.
3.3 Approximate cross validation
The magnitude of the error is not the only challenge for a brute-force approach. A practical issue is the sheer amount of time and/or computational power used to run a brute-force leave-one-cell-out approach. For example, in the first iteration of the previously described simulation study, the posterior inference had to be rerun $294$ times (once for each cell plus one for full data), and this example is in many ways a relatively small toy problem in terms of the size of the poststratification matrix. Many MRP examples involve poststratifying to an area level variable like postcode, which might have $100$ levels (Machalek et al., 2022). In this section, we apply an approximate LOO method based on Pareto-smoothed importance-sampling (PSIS; Vehtari et al., 2017, 2022).
PSIS-LOO (Vehtari et al., 2017, Eqn 10) estimates the sum expected log individual predictive density as
$\widehat{\operatorname{elpd}}_{\mathrm{PSIS-LOO}}=\sum_{i=1}^{n}\log\Bigg{(}\frac{\sum_{b=1}^{B}w_{i}^{b}\mbox{Pr}(y_{i}|\theta^{b})}{\sum_{b=1}^{B}w_{i}^{b}}\Bigg{)},$
where $w_{i}^{b}$ is the Pareto-smoothed importance weight for the $b^{th}$ posterior draw (denoted $s$ by Vehtari et al., 2017) and the $i^{th}$ individual. We modify this for our purposes by using other scoring metrics. We begin with substituting the error in each cell as before, with for shorthand we denote $\epsilon_{j}^{b}=\mbox{Pr}(Y_{i\in j}=1|y_{i\notin j},\textbf{x},\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j}=1|n_{j},y_{j})$. For each cell $j$ we can compute Pareto-smoothed weights $w_{j}^{b}$ for each posterior sample $b$ to estimate $\epsilon_{j}^{b}$ as outlined by Vehtari et al. (2017). Using these, we can approximate $\epsilon_{j}^{b}$ and substitute into (16) as
$\widehat{\operatorname{Error}}_{\mathrm{PSIS-LOCO,MRP}}^{2}=\Bigg{(}\frac{1}{N}\sum_{j=1}^{J}N_{j}\frac{\sum_{b=1}^{B}w_{j}^{b}\epsilon_{j}^{b}}{\sum_{b=1}^{B}w_{j}^{b}}\Bigg{)}^{2}.$ (18)
and thus estimate the squared error of the MRP estimate using the full model fit.
It is slightly more complicated to achieve this with the CRPS, but we begin by denoting the $g(y_{i\notin j},\textbf{x},\hat{\theta}^{b},\hat{\theta}^{\prime b})=\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{\prime b},\textbf{x})$ and $h(y_{i\notin j},\textbf{x},y_{j},n_{j},\hat{\theta}^{b})=\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|n_{j},y_{j})$. This means we can write (17) more concisely as
$\begin{split}\widehat{\operatorname{CRPS}}_{\mathrm{MRP}}(y_{i\notin j},y_{j},\textbf{x},\hat{\theta}^{b},\hat{\theta}^{\prime b})=\frac{1}{2}\frac{1}{B}\sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}g(y_{i\notin j},\textbf{x},\hat{\theta}^{b},\hat{\theta}^{\prime b})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}-\\ \frac{1}{B}\sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}h(y_{i\notin j},\textbf{x},y_{j},n_{j},\hat{\theta}^{b})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.\end{split}$
While it is possible to simply take an importance-weighted sum for the squared error, CRPS doesn’t have an amenable form. To apply the PSIS weights, which are the $b^{th}$ weight for the $j^{th}$ cell, we do importance resampling of the posterior draws for each cell using a stratified resampling algorithm (Kitagawa, 1996) using the $B$ PSIS weights for each cell. This procedure is implemented in the posterior R package (Bürkner et al., 2023). This produces $B$ resampled posterior draws $\hat{\theta}_{RW}$, which we permute to calculate PSIS reweighted $\hat{\theta}_{RW}^{\prime}$, and thus $g(y_{i\notin j},\textbf{x},\hat{\theta}_{RW}^{b},\hat{\theta}_{RW}^{\prime b})$ and $h(y_{i\notin j},\textbf{x},y_{j},n_{j},\hat{\theta}_{RW}^{b})$, to get
$\begin{split}\widehat{\operatorname{CRPS}}_{\mathrm{PSIS-LOCO,MRP}}(y_{i\notin j},y_{j},\textbf{x},\hat{\theta}^{b},\hat{\theta}^{\prime b})=\frac{1}{2}\sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}g(y_{i\notin j},\textbf{x},\hat{\theta}_{RW}^{b},\hat{\theta}_{RW}^{\prime b})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}-\\ \sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}h(y_{i\notin j},\textbf{x},y_{j},n_{j},\hat{\theta}_{RW}^{b})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.\end{split}$ (19)
Figure 5 shows that the approximate PSIS-LOCO scores are approximately equivalent to the brute-force LOCO scores.
Figure 5: Comparison of the fast Pareto-smoothed importance sampling approximation ($y$-axis) against the brute-force MRP LOCO estimate ($x$-axis). Colour of the point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents an identity line. The strong relationship indicates that this approximation is highly efficient, with approximate methods closely following brute-force methods, with a small amount of noise.
Through these simulations, we demonstrate the efficacy of our proposed cellwise scoring in a leave-one-cell-out approach. Whilst the model ordering is well preserved (which is desirable as other approaches do not preserve this), the scores as estimated with the sample are more favourable to the quality of the estimate than they should be. Using a leave-one-cell-out approach only marginally improves on this, but can be well approximated with a fast approximation such as PSIS-LOCO. Although it is frustrating to underestimate to this degree, we feel the preservation of model ordering still makes this a useful tool, albeit one where the scale must be interpreted carefully.
When we followed the traditional approach of comparing the mean of cell scores to the true mean of cell scores, the leave-one-cell-out approach appeared to slightly over estimate the magnitude of error (see supplementary), suggesting that the observed underestimation is a feature of our particular score decomposition.
4 Differentiating between population and subpopulation goodness
One of the benefits of MRP is that the same model can be used for both population and subpopulation estimates. To demonstrate how, we follow Kuh et al. (2022) to use set notation to describe the set of cells that form the subpopulation as $\mathbb{S}$. The total number of cells in this set is $S$ and any particular cell as $s$. In the case of the full population, $\mathbb{S}$ is all cells in the population, $S=J$ and $s=j$. Thus we can express a population weighted estimate for any subpopulation (small area) in the population as
$\operatorname{E}_{\mathrm{S}}(Y|y,\textbf{x},\hat{\theta})=\frac{\sum_{s=1}^{S}N_{s}\mbox{Pr}(Y_{i\in s}=1|y,\textbf{x},\hat{\theta})}{\sum_{s=1}^{S}N_{s}}.$ (20)
One benefit of the leave-one-cell-out estimation technique described in the previous section is that it is also constructed within each cell. This means that we can apply both the CRPS score and squared-error score described previously by focusing on the set of cells that describe a particular subpopulation. In the previous section we demonstrated the accordance of the PSIS-LOCO approximation to brute-force LOCO, and so in this section we use only the PSIS-LOCO for computational efficacy.
In this study we add two more models to our comparison set to add great distinction for the $X_{1}$ and $X_{3}$ subpopulation estimates:
  • $X_{1}\mathrm{only}$: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{1})}_{j}),$
  • $X_{3}\mathrm{only}$: $\mbox{Pr}(y_{j}=1|n_{j})=\operatorname{logit}^{-1}(\beta_{0}+\alpha^{(X_{3})}_{j}),$
We can modify (18) to
$\widehat{\operatorname{Error}}_{\mathrm{S,PSIS-LOCO,MRP}}^{2}=\Bigg{(}\frac{1}{\sum_{s=1}^{S}N_{s}}\sum_{s=1}^{S}N_{s}\frac{\sum_{b=1}^{B}w_{s}^{b}\epsilon_{s}^{b}}{\sum_{b=1}^{B}w_{s}^{b}}\Bigg{)}^{2}$ (21)
by simply limiting to the relevant subpopulation rather than the full population. Similarly we can estimate the CRPS for the subpopulaton by modifying (19) to similarly focus on the relevant subpopulation:
$\begin{split}\widehat{\operatorname{CRPS}}_{\mathrm{S,MRP}}(y,n,\hat{\theta},\mathbf{x})=\frac{1}{2}\sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{s=1}^{S}N_{s}g(y_{i\notin s},\textbf{x},\hat{\theta}_{RW}^{b},\hat{\theta}_{RW}^{\prime b})}{\sum_{s=1}^{S}N_{s}}\Bigg{\rvert}\Bigg{)}-\\ \sum_{b=1}^{B}\Bigg{(}\Bigg{\lvert}\frac{\sum_{s=1}^{S}N_{s}h(y_{i\notin s},\textbf{x},y_{s},n_{s},\hat{\theta}_{RW}^{b})}{\sum_{s=1}^{S}N_{s}}\Bigg{\rvert}\Bigg{)},\end{split}$ (22)
using the same importance resampling technique as described previously.
It is common for multiple subpopulations to be estimated, most often all levels of a variable (e.g., geographic regions). This means there is interest in scoring the goodness of estimates for all levels of a particular variable rather than one specific level. To describe this we use $l$ to denote a particular level of the $k^{th}$ variable, with the total number of levels in this variable $L$. Where we it is necessary to identify a specific variable we denote with a superscript $(k)$ but otherwise assume this is clear from the context for simplicity. For a particular score, it is possible to plot and consider $\operatorname{Score}_{l}$, for every $l$, but it is also possible to consider the average of scores for every level:
$\operatorname{Score}^{(k)}=\frac{\sum_{l=1}^{L^{(k)}}\operatorname{Score}_{l}}{L^{(k)}},$ (23)
which in the case of the score being squared error, corresponds to the mean squared error over levels of $k$.
Figure 6 shows a comparison between the mean over levels of the estimated PSIS-LOCO score for each variable and the true mean over levels. Focusing on the two variables that are strongly predictive of the outcome ($X_{2}$ and $X_{4}$, final two columns) we can cleanly distinguish between models with a low mean true score and models with a high mean true score. This was a demonstrated challenge in Kuh et al. (2022) that these new metrics overcome. Of particular interest for the practitioner is that the models with the lowest true score differ for the different variables, which again reinforces the concept proposed by Kuh et al. (2022) that the best model differs based on the overall aim.
Turning our attention to the variables that are less predictive of the outcome ($X_{1}$ and $X_{3}$, first two columns), we see less clear delineation between good and bad models. However, we still retain the respective ordering despite the noisiness of estimates. The respective score of each level could be plotted individually; see supplementary materials.
Figure 6: Comparison the mean of true scores across levels ($y$-axis) against the mean of estimated scores ($x$-axis, alternating vertical facets) for each level in each variable (horizontal and vertical facets). Colour of the point represents the different models fitted, whilst different points within this represent different simulation iterations. The black line represents an identity line. These results can be separated into two classes, those estimating levels of X2 and X4 (strongly predictive of outcome) and those estimating levels of X1 and X3 (weakly predictive of outcome).
5 Reference validation score
To use our scoring decomposition we require at least one observation in each cell. However, in many real-world cases not all cells will be observed in the sample. To overcome this, we follow the approach of Vehtari & Ojanen (2012) and substitute the observed cellwise probability with an estimated probability using a reference model denoted $M_{*}$. In this section we first consider a full reference validation score where we make this substitution for every cell in Section 5.1. This is so we can compare to previous approaches. Then, in Section 5.2 we consider the case where not all cells are observed. In this section we first extend the PSIS LOCO reference validation to this scenario. The we propose a method we call LOCO combined validation, which is a combination of LOCO cross validation and LOCO reference validation. Finally we propose a method to validate the chosen reference model.
5.1 All cells observed
The reference validation technique follows the approach of Vehtari & Ojanen (2012). We denote the reference model as $M_{*}$, and compare to a candidate model, $M_{c}$, with the overall aim to evaluate if they perform similarly in estimating the population expectation. The reference model can be difficult to propose (as discussed by Vehtari & Ojanen, 2012), but in an MRP context there is often need to choose the smallest/simplest model. In this paper the model with $X_{4}$ could be used almost interchangeably to the model with all predictors. This is important as the demographic variables need to be estimated in the population, which can be imprecise or require additional modelling (e.g., Gellar et al., 2023).
Similarly, although a model with more complex priors might reduce individual error (e.g., Gao et al., 2021), if the target estimand is at the population level, the additional model and computational complexity might not be justified if the estimate is similarly good with a simpler model.
We can apply a reference validation approach by substituting the reference model estimate for the population truth in (6) and (10). In Equation (30) of Vehtari & Ojanen (2012), the expected loss function under the actual belief model $M_{*}$ is decomposed in the variance for a predicted value given $M_{*}$ and the squared difference in expectations between the reference model $M_{*}$ and the proposed model $M_{c}$. As mentioned by Vehtari & Ojanen (2012) the model minimising the squared loss is that which the predicted expectation is closest to that with the reference model. Incorporating this into (6) in an MRP context, we would minimise
$\operatorname{Error}_{\mathrm{MRP}}(M_{c},M_{*})^{2}=(E_{P}(Y|M_{c},y,\textbf{x})-E_{P}(Y|M_{*},y,\textbf{x}))^{2},$ (24)
which can be expanded in a similar form as in previous sections to
$\operatorname{Error}_{\mathrm{MRP}}(M_{c},M_{*})^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}=1|M_{c},y,\textbf{x})-\mbox{Pr}(Y_{i\in j}=1|M_{*},y,\textbf{x}))}{\sum_{j=1}^{J}N_{j}}\Bigg{)}^{2}.$ (25)
Similarly we use Lemma (17) of Székely & Rizzo (2005) to measure the difference between the predictive distribution of $M_{*}$ and $M_{c}$. This lemma is also used by Gneiting & Raftery (2007) when discussing CRPS:
[…]
where G(t) is the CDF for the candidate model and F(t) is the CDF for the reference model and so we use $\mbox{Pr}(Y|M_{*},y,\textbf{x})$ as Y and $\mbox{Pr}(Y|M_{c},y,\textbf{x})$. Subbing these in and simplifying as in previous sections, we get
$\frac{1}{2}\sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{c},y,\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{c}^{\prime},y,\textbf{x}\bigg{)}\Bigg{|}\,+\\ \frac{1}{2}\sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{*},y,\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{*}^{\prime},y,\textbf{x}\bigg{)}\Bigg{|}\,-\\ \sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{c},y,\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{*},y,\textbf{x}\bigg{)}\Bigg{|}\,.$ (27)
Where we have observed cells, it is important to use the LOCO estimate of $\mbox{Pr}(y_{i\in j}=1|M_{c},y,\textbf{x})$, $\mbox{Pr}(y_{i\in j}=1|M_{c},y_{\notin j},\textbf{x})$ to avoid overfitting and more accurately capture the predictive power for unseen observations. Specifically for squared error, we have
$\operatorname{Error}_{\mathrm{MRP},\mathrm{LOCO}}(M_{c},M_{*})^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}\left(Y_{i\in j}=1|M_{c},y_{i\notin j},\textbf{x}\right)-\mbox{Pr}\left(Y_{i\in j}=1|M_{*},y_{i\notin j},\textbf{x}\right))}{N}\Bigg{)}^{2},$ (28)
and for CRPS we have
$\!\!\!\!\!\!\!\!\!\!\!\!\operatorname{CRPS}_{\mathrm{MRP},\mathrm{LOCO}}(M_{c},M_{*})=$
$\sum_{b=1}^{B}\Bigg{(}\frac{1}{2B}\ \Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{c},y_{i\notin j},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{c}^{\prime},y_{i\notin j},\textbf{x}\bigg{)}\Bigg{|}\,+$
$\frac{1}{2}\ \Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{*},y_{i\notin j},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{*}^{\prime},y_{i\notin j},\textbf{x}\bigg{)}\Bigg{|}\,-$
$\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\mbox{Pr}(y_{i\in j}=1|M_{c},y_{i\notin j},\textbf{x})-\mbox{Pr}(y_{i\in j}=1|M_{*},y_{i\notin j},\textbf{x}\bigg{)}\Bigg{|}\,\Bigg{)}.$
The PSIS approximation of the LOCO reference model estimate for each cell can be substituted. For the squared error score, we simply use an importance-weighted estimate for each cell:
$\operatorname{Error}_{\mathrm{MRP},\mathrm{LOCO},\mathrm{PSIS}}(M_{c},M_{*})^{2}=$ (29)
$\!\!\!\!\!\!\!\!\!\!\Bigg{(}\frac{1}{N}\sum_{j=1}^{J}N_{j}\bigg{(}\frac{\sum_{b=1}^{B}w_{M_{c},j}^{b}\mbox{Pr}(Y_{i\in j}=1|M_{c},\hat{\theta}^{b}_{M_{c}}y_{i\notin j},\textbf{x})}{\sum_{b=1}^{B}w_{M_{c},j}^{b}}-\frac{\sum_{b=1}^{B}w_{M_{*},j}^{b}\mbox{Pr}(Y_{i\in j}=1|M_{*},\hat{\theta}^{b}_{M_{*}}y_{i\notin j},\textbf{x})}{\sum_{b=1}^{B}w_{M_{*},j}^{b}}\bigg{)}\Bigg{)}^{2}.$
Figure 7: Comparison of the candidate model LOCO cross validation score (when estimated using Pareto-smoothed importance sampling, $y$-axis) against the LOCO reference validation approach ($x$-axis). Colour of point represents the different candidate models, whilst different points within this represent different simulation iterations. The black line represents an identity line. The strong relationship indicates that when the reference model is well chosen, this closely approximates the PSIS-LOCO score.
For the LOCO reference validation CRPS score we can use the importance-resampled posteriors for reference model $M_{*}$, $\hat{\theta}_{RW,M_{*}}$ and candidate model $M_{c}$, $\hat{\theta}_{RW,M_{c}}$ We then modify as follows: $g_{\mathrm{REF}}(y_{i\notin j},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{c}}^{\prime b})=\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}_{RW,M_{c}}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}_{RW,M_{c}}^{\prime b},\textbf{x})$ and $h_{\mathrm{REF}}(y_{i\notin j},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{*}}^{b})=\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}_{RW,M_{c}}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}|y_{i\notin j},\hat{\theta}_{RW,M_{*}}^{b},\textbf{x})$.
$\operatorname{CRPS}_{\mathrm{MRP},\mathrm{LOCO}}(M_{c},M_{*})=\frac{1}{2B}\Bigg{(}\sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}g_{\mathrm{REF}}(y_{i\notin j},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{c}}^{\prime b})\bigg{)}\Bigg{|}\,+$
$\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}g_{\mathrm{REF}}(y_{i\notin j},\textbf{x},\hat{\theta}_{RW,M_{*}}^{b},\hat{\theta}_{RW,M_{*}}^{\prime b})\Bigg{|}\,-\Bigg{|}\frac{1}{N}\sum_{j=1}^{J}N_{j}h_{\mathrm{REF}}(y_{i\notin j},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{*}}^{b})\Bigg{|}\,\Bigg{)}.$
To demonstrate the efficacy of the PSIS LOCO reference validation approach, we let the full model with all covariates be $M_{*}$ and the candidate models be the model without the precision variable, the model without the bias variable, and the model without either. For each we use a PSIS approximation of the predictive error of cell $j$ and apply this to our previous example with all cells observed. Figure 7 demonstrates that the PSIS LOCO reference validation approach and the PSIS LOCO cross validation approach described in Section 3 produce similar estimates.
These results are conditional on choosing an appropriate reference model. The particular example presented is a best case scenario as the reference model aligns closely with the data generating process. In the following section we discuss a technique to validate the reference model and a method for combining cell scores for observed and unobserved cells.
5.2 Not all cells observed
In this section we discuss the (common) case when not all cells are observed. We contribute three components. Firstly we detail how to use the reference validation approach in this case. Secondly, we propose a tool to validate the reference model. Thirdly, we propose a technique to combine the cell reference model score and LOCO cross validation model score.
5.2.1 Reference validation score
We begin by decomposing the population into two sets. The first set, $\mathbb{S}_{\mathrm{obs}}$, is the set of cells where at least one observation in each is observed in the sample. The total number of cells in this set is $S_{\mathrm{obs}}$ and any particular cell in this set is $s_{\mathrm{obs}}$. The second set, $\mathbb{S}_{\mathrm{unobs}}$, is the set of cells that do not have any observations in the sample. The total number of cells in this set is $S_{\mathrm{unobs}}$ and any particular cell in this set is $s_{\mathrm{unobs}}$. For clarity note that $J=S_{\mathrm{obs}}+S_{\mathrm{unobs}}$.
To compose the reference validation score, we note that when cells are observed we should use a LOCO approach to assess predictive power. When cells are not observed, we can simply use the difference in prediction between model $M_{c}$ and model $M_{*}$. This means we can write a score that combines PSIS approximated LOCO predictions with predictions for unobserved cells by modifying (29). For simplicity we denote
$\mbox{Pr}(Y_{i\in j}=1|\hat{\theta}^{b}_{M_{c}}y_{i\notin j},\textbf{x},w_{M_{c},j})=\frac{\sum_{b=1}^{B}w_{M_{c},j}^{b}\mbox{Pr}(Y_{i\in j}=1|\hat{\theta}^{b}_{M_{c}}y_{i\notin j},\textbf{x})}{\sum_{b=1}^{B}w_{M_{c},j}^{b}}$
and our modified equation as
$\operatorname{Error}_{\mathrm{REF}}(M_{c},M_{*})^{2}=$
$\!\!\!\!\!\!\!\!\!\!\!\!\Bigg{(}\frac{1}{N}\Bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}\!\!N_{s_{\mathrm{obs}}}\bigg{(}\mbox{Pr}(Y_{i\in s_{\mathrm{obs}}}=1|\hat{\theta}^{b}_{M_{c}}y_{i\notin s_{\mathrm{obs}}},\textbf{x},w_{M_{c},s_{\mathrm{obs}}})-\mbox{Pr}(Y_{i\in s_{\mathrm{obs}}}=1|\hat{\theta}^{b}_{M_{*}}y_{i\notin s_{\mathrm{obs}}},\textbf{x},w_{M_{*},s_{\mathrm{obs}}})\bigg{)}+$
$\sum_{s_{unobs}=1}^{S_{unobs}}N_{s_{unobs}}\bigg{(}\mbox{Pr}(Y_{i\in s_{\mathrm{unobs}}}=1|\hat{\theta}_{M_{c}},y,\textbf{x})-\mbox{Pr}(Y_{i\in s_{\mathrm{unobs}}}=1|\hat{\theta}_{M_{*}},y,\textbf{x})\Bigg{)}\Bigg{)}^{2}.$ (30)
Each component of CRPS can similarly be decomposed into an approximate LOCO value for the cells with observations and a prediction the cells with no sample observations. We denote $g^{\prime\prime}_{\mathrm{REF}}(y_{j+1},\textbf{x},y,\hat{\theta}_{M_{c}}^{b},\hat{\theta}_{M_{c}}^{\prime b})=\mbox{Pr}(Y_{i\in j+1}|y,\hat{\theta}_{M_{c}}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j+1}|y,\hat{\theta}_{M_{c}}^{\prime b},\textbf{x})$ and $h^{\prime\prime}_{\mathrm{REF}}(y,\textbf{x},\hat{\theta}_{M_{c}}^{b},\hat{\theta}_{M_{*}}^{b})=\mbox{Pr}(Y_{i\in j+1}|y,\hat{\theta}_{M_{c}}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j+1}|y,\hat{\theta}_{M_{*}}^{b},\textbf{x})$ as the corresponding $g(.)$ and $h(.)$ for the cells with no sample observations.
$\operatorname{CRPS}_{\mathrm{MRP},\mathrm{LOCO}}(M_{c},M_{*})=$
$\ \ \ \ \frac{1}{B}\ \Bigg{(}\frac{1}{2}\sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}g_{\mathrm{REF}}(y_{i\notin s_{\mathrm{obs}}},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{c}}^{\prime b})+$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}N_{s_{\mathrm{unobs}}}g^{\prime\prime}_{\mathrm{REF}}(y_{s_{\mathrm{unobs}}},\textbf{x},y,\hat{\theta}_{M_{*}}^{b},\hat{\theta}_{M_{*}}^{\prime b})\bigg{)}\Bigg{|}\,+$
$\ \ \ \ \ \ \ \ +\ \frac{1}{2}\sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}g_{\mathrm{REF}}(y_{i\notin s_{\mathrm{obs}}},\textbf{x},\hat{\theta}_{RW,M_{*}}^{b},\hat{\theta}_{RW,M_{*}}^{\prime b})+$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}N_{s_{\mathrm{unobs}}}g^{\prime\prime}_{\mathrm{REF}}(y_{s_{\mathrm{unobs}}},\textbf{x},y,\hat{\theta}_{M_{c}}^{b},\hat{\theta}_{M_{c}}^{\prime b})\bigg{)}\Bigg{|}\,+$
$\ \ \ \ \ \ \ \ +\ \sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}h_{\mathrm{REF}}(y_{i\notin s_{\mathrm{obs}}=1},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{*}}^{b})+$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}N_{s_{\mathrm{unobs}}}h_{\mathrm{REF}}(y_{i\notin s_{\mathrm{unobs}}=1},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{*}}^{b})\bigg{)}\Bigg{|}\,\Bigg{)}.$ (31)
5.2.2 Combined validation score
The challenge with the method described in Section 5.2.1 is that it doesn’t use the information that we have observed in the sample to test model goodness. In this section we utilise this information in two ways. Firstly we propose a technique to provide partial validation for the reference model. Secondly, we propose a modification of (30) and (31) to use LOCO reference validation when the cells are not observed and LOCO cross validation when they are.
Firstly we focus on the portion of the population that is observed. We subset to this section of the population and calculate the PSIS LOCO reference validation score and the PSIS LOCO cross validation score by treating the observed cells as a subpopulation; see Section 4. By comparing the resultant model ordering we can identify whether the two approaches suggest similar ordering for the subpopulation where the sample has an observation in each cell.
To demonstrate this, we extend the simulations in Section 2 to relax the assumption that every cell needs to be observed in the sample. Instead we simply constrain that every level of a variable needs to be observed in the sample when taking a sample from the population. With this constraint relaxed, our simulated samples contained $28.9\%$ of the cells in the population (min $24.9\%$ and max $33.1\%$). To begin with we focus on the portion of the population poststratification matrix observed in the sample.
Figure 8 plots the PSIS LOCO cross validation scores scores against the reference model scores. The first column uses the full model as the reference model, whilst the second column uses the precision model as the reference model. When the reference model is chosen appropriately (the full model) the two scores align, but when an inappropriate reference model is chosen (the precision model), they do not.
Figure 8: Comparison of the PSIS LOCO cross validation score ($x$-axis) against the PSIS LOCO reference validation score ($y$-axis). Colour of point represents the different candidate models, whilst different points within this represent different simulation iterations. The top row represents the CRPS scores when the reference model is the full model (left panel) and the precision model (right panel). The bottom represents the same for the squared error.
However, this only evaluates part of the population and should not be used to score the models. Instead we propose an approach that combines the reference validation score and the cross validation score. In the reference validation approach we don’t leave out anything when we don’t have any observation in that cell, and in the combined approach the reference model approach is used only when there is no observation in that cell. For squared error we modify (30) to
$\operatorname{Error}_{\mathrm{COMB}}(M_{c},M_{*})^{2}=\Bigg{(}\frac{1}{N}\Bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}\bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}\frac{\sum_{b=1}^{B}w_{s_{\mathrm{obs}}}^{b}\epsilon_{\mathrm{obs}}^{b}}{\sum_{b=1}^{B}w_{s_{\mathrm{obs}}}^{b}}\bigg{)}\ +$
$\ \ \ \ \sum_{s_{unobs}=1}^{S_{unobs}}\!\!N_{s_{unobs}}\!\bigg{(}\mbox{Pr}\left(Y_{i\in s_{\mathrm{unobs}}}=1\left|\hat{\theta}_{M_{c}},y,\textbf{x}\right.\right)-\mbox{Pr}\left(Y_{i\in s_{\mathrm{unobs}}}=1\left|\hat{\theta}_{M_{*}},y,\textbf{x}\right.\!\right)\!\Bigg{)}\!\Bigg{)}^{2}\!\!,$ (32)
where $w_{\mathrm{obs}}^{b}$ and $\epsilon_{\mathrm{obs}}^{b}$ are as defined in Section 3.3. Similarly we take (31) and modify to use $g(.)$ and $h(.)$ as defined in Section 3.3 where the cell is observed:
$\ \ \ \ \ \ \ \ \ \ \ \ \ \operatorname{CRPS}_{\mathrm{MRP},\mathrm{LOCO}}(M_{c},M_{*})=$
$\!\!\!\!\!\!\!\!\frac{1}{N}\Bigg{(}\sum_{b=1}^{B}\frac{1}{2B}\ \Bigg{|}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}g(y_{i\notin s_{\mathrm{obs}}},\textbf{x},\hat{\theta}_{M_{c}}^{b},\hat{\theta}_{M_{c}}^{\prime b})+\!\!\!\!\sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}\!\!N_{s_{\mathrm{unobs}}}g^{\prime\prime}_{\mathrm{REF}}(y_{s_{\mathrm{unobs}}},\textbf{x},y,\hat{\theta}_{M_{*}}^{b},\hat{\theta}_{M_{*}}^{\prime b})\Bigg{|}\,+$
$\frac{1}{2}\ \Bigg{|}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}g(y_{i\notin s_{\mathrm{obs}}},\textbf{x},\hat{\theta}_{M_{*}}^{b},\hat{\theta}_{M_{*}}^{\prime b})+\!\!\!\!\sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}\!\!N_{s_{\mathrm{unobs}}}g^{\prime\prime}_{\mathrm{REF}}(y_{s_{\mathrm{unobs}}},\textbf{x},y,\hat{\theta}_{M_{c}}^{b},\hat{\theta}_{M_{c}}^{\prime b})\Bigg{|}\,+$
$\Bigg{|}\sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}\!\!N_{s_{\mathrm{unobs}}}h_{\mathrm{REF}}(y_{i\notin s_{\mathrm{unobs}}=1},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{*}}^{b})\Bigg{|}\,\Bigg{)}.$ (33)
We lose the observed cell component of the final term as there is no distribution for these estimates in the observed cell case.
In Figure 9 we demonstrate the efficacy of this approach against the true model score. Similar to previous results, the magnitude of the score is underestimated for both scores but ordering is maintained.
Figure 9: Comparison of the true score for the estimate ($y$-axis) against the combined validation approach ($x$-axis). Colour of point represents the different candidate models, whilst different points within this represent different simulation iterations. Points below the black line represent underestimation of the magnitude for CRPS and above the identity line represent underestimation of the magnitude for squared error.
6 Conclusion
In this manuscript we propose adaptions to the squared error and continuous ranked probability score (CRPS) for use in multilevel regression and poststratification (MRP) settings. We demonstrate that this adaption correctly recovers model ordering at the population and subpopulations. We also provide an approximate leave-one-cell-out approach for fast estimation. We propose a reference validation approach and a combined validation approach for use when not all cells are observed in the sample. Together we feel this provides encouraging progress towards MRP model validation and consider this a considerable improvement on previous model validation attempts (which do not retain correct ordering).
This work provides evidence that failure to correctly recover model ordering is based in the sum of individual scores. However, our method does consistently underestimate the magnitude of the loss, which should be further investigated. Until a solution has been identified, the magnitude of loss should not be directly used as an approximation of loss of the true error.
In this manuscript we have focused primarily on the simulation design proposed by (Kuh et al., 2022). However, future work should consider different sampling conditions and model comparisons to ensure generalisability. This example is severe and may not be illustrative of real world applications. It could be useful to consider the scoring of models with different priors.